OpenStudy (anonymous):

F=9/5C+32 Find C when F = 53.6.

OpenStudy (anonymous):

You have F, which is 53.6, so to find C we will plug in 53.6 for F and evaluate for C: \[53.6=(9/5)C + 32\]You now need to get C by itself. You can start by subtracting 32 from both sides: \[21.6=(9/5)C\]Now, all you need to do is multiply both sides by 5/9 and you'll have isolated C.

OpenStudy (anonymous):

f=9/5c+32 now f=53.6 53.6=9/5c+32 53.6-32=9/5c 21.6=9/5c now cross multiply 5c=9/53.6 to make it easy multiply and divide it by 10 5c=90/536 5c=0.167 c=0.33 you can take 90/167 as many desimals you want depending upon answer